OpenStudy (anonymous):

find w... 15/w = w/15-w

OpenStudy (anonymous):

\[\frac{ 15 }{ w }=\frac{ w }{ 15 }-w\]

OpenStudy (anonymous):

get a common denominator on the right so you can get one fraction

OpenStudy (anonymous):

\[ \frac{ 15 }{ w } = \frac{ w }{ 15-w }\]

OpenStudy (anonymous):

Sry, I just realized I was using the wrong equation

OpenStudy (anonymous):

cross multiply your equation

OpenStudy (anonymous):

225-15w=w^2

OpenStudy (anonymous):

set it equal to zero

OpenStudy (anonymous):

0 = w^2 + 15w -225

OpenStudy (anonymous):

Use the quadratic formula to solve for w

OpenStudy (anonymous):

When you have the form\[ax^2+bx+c=0\]Quadratic formula is\[x=-b \pm \frac{ \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

\[w=-15\pm \frac{ \sqrt{15^2-4*1*-225} }{ 2*1 }\]

OpenStudy (anonymous):

sry for the confusion. Do the math above and you should have your answer. There will be 2 answers

OpenStudy (anonymous):

one was 9.3, i guess thats the more reasonable answer cause the other was negative